OpenStudy (anonymous):

Let log P/N=8 and log M/N=5, What is the relationship between P and M? a. p=100M b. p=0.001M c. p=3M d. p=1,000M

OpenStudy (anonymous):

probably a snappier way to do this, but you can write each one in equivalent exponential form \[\frac{P}{N}=10^8, \frac{M}{N}=10^5\] so \[P=10^8N,M =10^5N\] and therefore \[\frac{P}{M}=\frac{10^8N}{10^5N}=10^3\]

OpenStudy (anonymous):

So 1000.